OpenStudy (anonymous):

Find the exact value of the expression. Do not use a calculator. If tan θ = 9, find the exact value of cot (pi/2 - 0).

OpenStudy (anonymous):

@JULIANBMXR16

OpenStudy (anonymous):

hold on a minute.

OpenStudy (anonymous):

\[\large \cos(\frac{\pi}{2}-\theta)=\tan(\theta)\] this is one of the "cofunction" identities

OpenStudy (anonymous):

I tried this and got an answer that wasn't one of my choices, I don't understand.

OpenStudy (anonymous):

A.8 B.10 C.1/9 D.9